# Mamba State-Space Models Can Be Strong Downstream Learners

### v1

@@ 9 Acknowledgements
Appendix A Mamba stable dynamics proof
Recall the state-space parameters and equations for the MambaBlock; $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t},\bm{\Delta}_{t}\in\mathbb{R}{}^{d\times d}$ for $t\in\{1,\dots,n\}=[n]$. Given an input sequence $\mathbf{u}_{1},\dots,\mathbf{u}_{n}\in\mathbb{R}{}^{d}$, the following linear mapping through latent states $\bm{x}_{1},\dots,\bm{x}_{n}\in\mathbb{R}{}^{d}$ is used to produce the output $\mathbf{y}_{1},\dots,\mathbf{y}_{n}\in\mathbb{R}{}^{d}$:
@@ Appendix A Mamba stable dynamics proof
Appendix B Proof of weight-tying using LoRA in the MambaBlock
@@
$\tilde{\mathbf{W}}=$ $\mathbf{W}+\mathbf{W}^{\prime}$
$=$ $\mathbf{W}+\mathbf{U}\mathbf{V},$
Recall the form of $\mathbf{W}$:
$=$ $\mathbf{W}[t-1,j]+\sum_{k=0}^{r-1}\mathbf{U}[t-1,k]\mathbf{V}[k,j].$
Appendix C Experimental Details
All model checkpoints were evaluated on all benchmarks and few-shot settings using the LM evaluation harness from Eleuther AI [16], version 0.4.2. Pythia and Mamba Huggingface checkpoints were used for all inference and fine-tuning experiments, e.g., EleutherAI/pythia-160m and state-spaces/mamba-130m-hf for the smallest respective models. All fine-tuning experiments were run using package versions Transformers 4.40.0.dev0, Accelerate 0.28.0, TRL 0.8.1, PyTorch 2.2.1+cu121, and PEFT 0.10.0.
@@ Appendix C Experimental Details
Outside of the OpenHermes hyperparameter search, all Alpaca and OpenHermes fine-tuning experiments used the following training recipe (adapted from [50]): AdamW_torch optimizer, cosine annealing schedule, no gradient accumulation, maximum norm of 1.0 for gradient clipping, and no warmup steps. Training epochs used for all Alpaca and OpenHermes experiments were three and one, respectively. For both Pythia and Mamba models, the learning rate and LoRA dimension $r$ were scaled to improve performance of smaller models (per-model values listed in Table 3).
@@
All experiments in Tables 1 and  2, Figures 1 and  2(b) were run using a single-GPU Nvidia A10G (24 GB total memory). For Pythia and Mamba ALL LoRA experiments in Figure 3, all experiments were run on an A10G, except for Mamba 2.8B, which exceeded A10G memory capacity and was run on an Nvidia H100 (80 GB total memory).
@@
Mamba size Pythia size learning rate LoRA $r$
130M 160M 1.0e-5 8
370M 410M 5.0e-5 16
790M 1B 1.0e-6 32
1.4B 1.4B 5.0e-6 64
2.8B 2.8B 5.0e-7 128
For the hyperparameter search results in Figure 4, all experiments were run using 8 H100 GPUs. SLL LoRA was used for Mamba and Pythia models. The range of hyperparameter values was as follows:
  • learning rate $\in\{1e-7,2e-7,5e-7,1e-6,2e-6,5e-6,1e-5,2e-5,5e-5,1e-4,2e-4,5e-4,1e-3,2e-3,5e-3\}$
  • LoRA dimension $r\in\{16,32,64,128,256\}$
  • warmup steps $\in\{0,1000,2000\}$
All other hyperparameters followed previous experiments.

### v2

@@ 7 Acknowledgments
Appendix A Proof of weight-tying using LoRA in the MambaBlock
Due to the low-level nature of Mamba’s prefix scan optimizations (discussed in Section 3), standard use of LoRA adapters is made difficult within Mamba’s SSM-layer. E.g., while $B_{t},C_{t}$ and $\Delta_{t}$ are conceptually PyTorch linear layers, their bundling in a contiguous memory block and careful manipulation makes appending a LoRA adapter on any of these individual matrices non-trivial (particularly, while respecting the highly specialized layout of each LoRA adapters targeted layer). However, we note that the overall design of the MambaBlock’s hardware optimizations may be leveraged to both efficiently learn the parameter-space for the majority of time-varying parameters (thus achieving PEFT) and regularize parameters during training (thus improving fine-tuning generalization).
Theorem 1.
Proof.
$\tilde{\mathbf{W}}=$ $\mathbf{W}+\mathbf{W}^{\prime}$
$=$ $\mathbf{W}+\mathbf{U}\mathbf{V},$
Recall the form of $\mathbf{W}$:
$=$ $\mathbf{W}[t-1,j]+\sum_{k=0}^{r-1}\mathbf{U}[t-1,k]\mathbf{V}[k,j].$
Appendix B Mamba stable dynamics proof
Recall the state-space parameters and equations for the MambaBlock; $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t},\bm{\Delta}_{t}\in\mathbb{R}{}^{d\times d}$ for $t\in\{1,\dots,T\}=[T]$. Given an input sequence $\mathbf{u}_{1},\dots,\mathbf{u}_{T}\in\mathbb{R}{}^{d}$, the following linear mapping through latent states $\bm{x}_{1},\dots,\bm{x}_{T}\in\mathbb{R}{}^{d}$ is used to produce the output $\mathbf{y}_{1},\dots,\mathbf{y}_{T}\in\mathbb{R}{}^{d}$:
@@ Appendix B Mamba stable dynamics proof
Appendix C Mamba stable outputs proof
@@
Assume $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially non-increasing over discrete-time. Then small changes to the output $\mathbf{y}_{t}$ are also exponentially non-increasing over discrete time.
@@
Recall that $\bm{x}_{T}=F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})$. Furthermore, recall from Equations 1 and  2, $\mathbf{y}_{t}=\mathbf{C}_{t}\bm{x}_{t}$, where $\mathbf{C}_{t}$ is diagonal.
Let
$\mathbf{y}_{T}$ $=G_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})=\mathbf{C}_{T}\bm{x}_{T}=\mathbf{C}_{T}F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1}).$
Consider $\varepsilon$-close inputs $(\bm{x}_{t-1},\mathbf{u}_{t})$ and $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$, and their respective outputs $\mathbf{y}_{t}$ and $\mathbf{y}_{t}^{\prime}$. Assume $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially non-increasing over discrete-time. That is, $\max|F_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-F_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
We thus have
$\max|\mathbf{y}_{t}-\mathbf{y}_{t}^{\prime}|$ $=\max|G_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-G_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|$
$=\max|\mathbf{C}_{N}F_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-\mathbf{C}_{N}F_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|$
$\propto\max|F_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-F_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|,$
where proportionality follows due to the diagonality of $\mathbf{C}_{N}$ and the vector-absolute value. Thus,
$\max|G_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-G_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|$ $\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$
Appendix D Experimental Details
All model checkpoints were evaluated on all benchmarks and few-shot settings using the LM evaluation harness from Eleuther AI [16], version 0.4.2. Pythia and Mamba Huggingface checkpoints were used for all inference and fine-tuning experiments, e.g., EleutherAI/pythia-160m and state-spaces/mamba-130m-hf for the smallest respective models. All fine-tuning experiments were run using package versions Transformers 4.40.0.dev0, Accelerate 0.28.0, TRL 0.8.1, PyTorch 2.2.1+cu121, and PEFT 0.10.0. All Mamba-2 models were run using mamba-ssm v2.2.2 using Huggingface checkpoints, e.g., state-spaces/mamba-130m for the smallest model.
@@ Appendix D Experimental Details
All Alpaca and OpenHermes fine-tuning experiments used the following training recipe (adapted from [50]): AdamW_torch optimizer, cosine annealing schedule, no gradient accumulation, maximum norm of 1.0 for gradient clipping, and no warmup steps. Training epochs used for all Alpaca and OpenHermes experiments were three and one, respectively. For both Pythia and Mamba models, the learning rate and LoRA dimension $r$ were scaled to improve performance of smaller models (per-model values listed in Table 2).
@@
All experiments were run using a single-GPU Nvidia A10G (24 GB total memory). For Pythia, Mamba, and Mamba-2 ALL LoRA experiments in Figure 3, all models followed the same training and PEFT recipes, save for Mamba-2 2.7B which required a LoRA $r$ dimension of 64 to fit in A10G memory.
@@
Mamba size Mamba-2 size Pythia size learning rate Mamba/Pythia LoRA $r$ Mamba-2 LoRA $r$
130M 130M 160M 1.0e-5 8 8
370M 370M 410M 5.0e-5 16 16
790M 780M 1B 1.0e-6 32 32
1.4B 1.3B 1.4B 5.0e-6 64 64
2.8B 2.7B 2.8B 5.0e-7 128 64
@@
Appendix E Mamba-2 mixed-precision inference pretrained LLM performance tables
Table 3: Mean full-precision (FP32) divergence in MMLU performance for mixed-precision inference. Divergence is averaged over {0, 1, 3, 5}-shot performance. Pretrained checkpoints are used for Mamba (M), Mamba-2 (M2), Pythia (P), OLMo [19] (O), and Phi-1.5 [33] (Phi) models.
Model M M2 P M M2 P M M2 P O M M2 P Phi M M2 P
Size 130m 130m 160m 370m 370m 410m   790m 780m 1b 1.4b 1.3b 1.4b 1.5b 2.8b 2.7b 2.8b
FP16 $\mu$ 0.03 0.07 0.35 0.05 0.05 0.06 0.21 0.12 0.05 0.04 0.04 0.04 0.07 0.03 0.15 0.26 0.12
BF16 $\mu$ 0.05 0.67 1.45 0.20 0.52 0.20 0.66 0.29 0.16 0.13 0.31 0.40 0.13 1.05 1.17 1.02 0.11
Appendix F Adapted PEFT-layer effects on Mamba-2 zero-shot performance
As described in Section 3.2, both Mamba and Mamba-2 store the majority of time-varying SSM parameters in a large memory buffer (instantiated as a PyTorch linear layer’s weight matrix), denoted as $\mathbf{W}$. For PEFT via LoRA, we ablate the impact of targeting $\mathbf{W}$, as well as the impact of Theorem 3, where time-varying variables accessed through $\mathbf{W}$ were shown to be regularized through weight tying.
Firstly, we instruction tune Mamba-2 models using LoRA by both adapting $\mathbf{W}$ and adapting all linear layers other than $\mathbf{W}$. Mamba-2 MambaBlocks contain two linear layers, thus the former targets only the $\mathbf{W}$ linear layer whereas the latter targets the other linear layer in each block. As shown in Table 4, PEFT with only $\mathbf{W}$ targeted within the MambaBlock near uniformly results in better performance than only targeting other layers, with the former outperforming the latter on 32 out of 35 natural language tasks. This makes intuitive sense since $\mathbf{W}$ represents the majority of Mamba’s time-varying parameters (as previously mentioned). Thus, while previous works [21, 11] have displayed the importance of Mamba’s time-varying parameters for pretraining performance, this verifies the importance of Mamba’s time-varying parameters for instruction tuning performance.
Table 4: Zero-shot performance for instruction tuned Mamba-2 models where: denotes the large memory buffer $\mathbf{W}$ containing the majority of temporal variables (described in Section 3.2) is targeted for LoRA adaptation, and denotes $\mathbf{W}$ is not adapted. The top-performance for each task per model is highlighted in bold .
Model $\mathbf{W}$ targeted? LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande
ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$
Mamba-2 130M 15.37 45.16 35.41 65.02 47.94 24.91 52.17
16.95 43.30 35.24 64.85 47.47 24.06 52.17
Mamba-2 370M 8.03 54.16 46.87 69.53 53.49 27.73 57.14
8.50 53.74 46.59 70.67 54.80 26.79 55.64
Mamba-2 780M 5.79 61.61 55.14 71.98 61.11 29.18 60.09
5.86 61.63 54.94 72.03 60.98 28.41 60.22
Mamba-2 1.3B 4.54 65.73 60.88 73.67 66.16 34.56 61.80
5.05 65.44 59.84 73.39 64.10 33.11 61.09
Mamba-2 2.7B 4.05 69.63 66.73 76.50 69.95 36.77 64.56
4.10 69.61 66.60 76.39 69.53 36.26 63.93
Next, we present a counterintuitive experiment to demonstrate the impact of Theorem 3. As before, we instruction tune Mamba-2 models using LoRA, focusing on the two largest model sizes. This time, however, we target all linear layers within Mamba-2’s MambaBlock and contrast this with targeting only $\mathbf{W}$ within the MambaBlock. Displayed in Table 5, we can see that adapting more SSM parameters does not necessarily lead to improved performance across the board. Rather, adapting only $\mathbf{W}$ for Mamba-2 outperforms adapting all linear layers on eight of the 14 natural language tasks.
The results in Table 5 thus present two cases: (a) regularization: only adapting parameters which lead to weight tying (by Theorem 3) and (b) increased learning capacity: adapting more parameters, at the cost of learning unregularized parameters. Through regularization via weight tying [42], the former leads to good generalization. In contrast, while adapting more parameters, (b) includes unregularized variables, leading to positive improvements on only a minority of tasks when compared to the fine-tuning of fewer, but regularized, parameters. We note that adapting all linear layers is often performed when fine-tuning attention-based LLMs using LoRA [12]. Thus, a similar regularization-vs-capacity tradeoff for Transformers LLMs may present itself by carefully studying LoRA’s affects when targeting specific linear layers in attention-based architectures.
Table 5: Zero-shot performance for instruction tuned Mamba-2 models. denotes only $\mathbf{W}$ was targeted for LoRA adaptation within the MambaBlock, and thus denotes both linear layers within each MambaBlock were adapted. The top-performance for each task per model is highlighted in bold.
Model Only $\mathbf{W}$ targeted? LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande
ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$
Mamba-2 1.3B 4.54 65.73 60.88 73.67 66.16 34.56 61.80
4.94 62.95 60.06 73.78 64.02 34.81 61.33
Mamba-2 2.7B 4.05 69.63 66.73 76.50 69.95 36.77 64.56
3.77 68.39 67.51 75.95 69.53 38.31 65.35